# UPS: Efficiently Building Foundation Models for PDE Solving via Cross-Modal Adaptation

### v2

@@ 1 Introduction
Figure 1: Effectiveness-efficiency trade-off. We plot the normalized error averaged across 7 different PDE families in PDEBench vs. the estimated compute for single-family models (FNO, ORCA) and unified models (MPP-B, DPOT-M, UPS-B). Compute is estimated by $\log_{10}$ (Num GPUs $\times$ Num Train Steps). UPS achieves the lowest error at the lowest cost among all unified models, striking a balance between effectiveness and efficiency. Beyond compute, UPS also uses 4x less data compared with existing unified models. See Table 1 for detailed results and Appendix A for plot details.
@@
Partial Differential Equations (PDEs) play a pivotal role in modeling and understanding real-world phenomena, such as fluid dynamics and heat transfer. Although there exists a rich body of classical PDE solvers [1, 2, 3] that are effective and mathematically proven, these solvers often incur substantial computational costs when used in practice, as they need to be re-run every time a coefficient or boundary condition changes. This motivates the development of neural operators [4, 5, 6], which use neural networks to approximate a solution map for a PDE family and can generalize to different initial/boundary conditions or coefficients. While existing neural operators [7, 8, 9] have demonstrated strong performance on various practical benchmarks [10, 11], most of them are designed to work with a single PDE family. Training a separate model for each PDE family remains costly.
@@
Several recent works, such as Subramanian et al. [12], MPP [13], and DPOT [^1] This work was done at the same time as ours. [14], have taken initial steps towards developing foundation models for PDE solving, learning unified operators that transfer across PDE families. These models are pretrained from scratch using extensive amounts of data and compute. For example, MPP is trained with over 80,000 PDE trajectories on 8 NVIDIA H100 GPUs for 200,000 steps. Despite the development costs, the resulting models are limited in generalization ability—all existing unified models focus on pretraining with 2D PDEs. Finally, as these models are developed using only PDE trajectories, they do not leverage meta information that could help distinguish between various PDE families, such as the name of the family and the coefficients.
@@
Figure 2: To adapt pretrained LLMs for PDE solving, UPS first transforms PDE of different dimensions, channels, and resolutions into a unified representation (left panel). Then, the data is processed with a unified architecture that integrates FNO layers, PDE metadata, and LLMs (right panel). The architecture is trained in two stages. In stage 1, we pretrain the embedding network using a joint loss that simultaneously optimizes (i) the distribution similarity between PDE features and text embeddings to align the modalities, and (ii) the prediction performance of extracted PDE features. In stage 2, we fine-tune the entire model on a dataset that combines multiple families of spatiotemporal PDEs with varying domain dimensions, initial/boundary conditions, and coefficients. UPS achieves remarkable results with significantly better sample-efficiency than existing methods.
@@
We present Unified PDE Solvers (UPS), which learns unified neural operators for complex time-dependent PDEs with improved efficiency and generalization ability. Unlike existing efforts that train models from scratch, we propose a novel method to adapt pretrained Large Language Models (LLMs) to PDE solving. This is inspired by the line of work that repurposes LLMs for scientific domains like mathematics [15], computational biology [16, 17, 18], and chemistry [19, 20]. These works not only show how LLMs utilize both text and non-text information to solve scientific problems and transfer effectively to unseen tasks, but also provide strong evidence that the general-purpose pretraining and inductive biases of LLMs could substantially reduce the sample complexity needed for adaptation.
@@
    We propose a unified data representation scheme to align PDEs with varying dimensions and physical quantities into the same feature space. Given the space and time discretization $\bm{u}=\{\bm{u}_{t}(\bm{x})\}_{t=0}^{T}$, where $\bm{x}\in\mathbb{R}^{d}$ is the spatial variable and $\bm{u}_{t}(\bm{x})$ is the state variable, UPS homogenizes $\bm{u}_{t}(\bm{x})$ from diverse PDEs into a shared “superspace" $\mathbb{R}^{N\times n^{d_{\max}}}$, where $d_{\max}$ is the maximum dimension of $\bm{x}$ among all PDEs considered, $N$ is the superset of the physical quantities, and $n$ is the resolution. This framework of embedding lower-dimensional PDEs in a higher dimension enables UPS to model cross-dimensional PDEs simultaneously and distinguishes us from all existing unified operators, which do not consider low dimensional PDEs in 1D.
@@
    We employ a unified network architecture to predict $\bm{u}_{t+1}(\bm{x})$ based on $\bm{u}_{t}(\bm{x})$. To leverage pretrained LLMs, we design a three-way architecture that consists of (i) a FNO [4] based embedding network to convert PDE data into resolution-invariant, sequential features; (ii) an LLM body to process the PDE features and the text embeddings of the PDE metadata; and (iii) a prediction network to generate the final output. Inspired by previous cross-modality adaptation work [16], we employ a two-stage align-then-refine process for model training. However, we improve the align stage by using a joint loss that adds feature extraction on top of alignment to pretrain the embedding network. We improve the refine stage by fine-tuning on a collection of PDE tasks rather than a single task. Our enhanced workflow outperforms naive transfer and previous cross-modal approaches, reducing both the data and compute needed for training.
By design, UPS can handle diverse PDE families, data dimensions, channels, and resolutions. More crucially, by warm-starting with pretrained LLM weights and applying explicit alignment, UPS strikes a balance between effectiveness and efficiency (Figure 1)—it achieves state-of-the-art performance across 9 datasets from PDEBench [10] (7 in-distribution, 2 out-of-distribution), using about 20,000 training data, a single A6000, 60,000 train steps, and under 100 GPU hours. This means that we achieve better results than existing unified models using 4 times less data and 26 times less compute.
Beyond prediction accuracy, we confirm that UPS preserves key properties of neural operators, such as grid- and resolution-invariance. We also show that UPS is compatible with a variety of LLM backbones, including RoBERTa [21], T5 [22], and CLIP [23], and demonstrates better performance when scaled to larger backbones. We believe that the model-agnostic design of UPS offers a systematic approach to harnessing the advancements in LLMs for PDE solving, and it takes a further step towards building generalized foundation models for more complex physical systems efficiently. Code is available at https://github.com/sjunhongshen/UnifiedPDESolvers.
@@ 2 Related Work
Recent years has seen a variety of neural-network-based methods for approximating PDE solutions. Hybrid solvers [24, 25, 26] apply classical solvers like finite element/volumn methods [2, 3] to a low-resolution grid and use neural networks to predict the correction terms. Others directly approximate the PDE solutions with neural networks [27, 28, 29, 30], using variational losses [31] or physical constraints defined by the PDE [28, 32]. Being mostly equation-specific, these methods can solve one PDE at a time. The learned models do not apply to other PDEs in the same family, let alone other families.
@@
A more general approach involves learning neural operators [6, 4, 33] which approximate an infinite-dimensional operator between two functional spaces. For time-dependent PDEs, a neural operator maps the current state of a PDE to the next state, with quantities like initial conditions provided as input. Neural operators can be implemented using any architecture. For example, Fourier neural operator (FNO) [4] uses convolution-based integral kernels evaluated in the Fourier space. Other works also use transformer models [34, 35, 8] or U-Net [7]. Learning neural operators enables solving an entire family of PDE and they can easily adapt to new parameterizations of a PDE without fine-tuning. However, the learned operators cannot extend to different PDE families.
@@
To facilitate operator transfer across PDE families, two recent works develop large pretrained models for multiple physical systems: Subramanian et al. [12] train FNOs on steady-state linear PDEs with periodic boundary conditions; McCabe et al. [13] design a new transformer architecture based on the axial attention [36] and train it using various 2D non-linear, time-dependent PDEs. While these methods show that a unified operator can outperform single-family operators, they are limited in two aspects:
  • Modeling-wise, existing unified methods consider mainly 2D PDEs for pretraining and evaluation. In contrast, UPS leverages a unified representation scheme to tackle both 1D and 2D PDEs. This method can be also extended to any $d$-dimensional systems in theory.
  • Efficiency-wise, existing methods pretrain large models from scratch and necessitate extensive GPU resources and pretraining data, which can be prohibitive to collect for high-dimensional complex PDEs. However, by adapting from pretrained LLMs and closing the modality gap between text and PDE efficiently, UPS achieves competitive results using 4x less data and 26x less compute.
@@
Beyond the aforementioned works, DPOT [14] was developed concurrently with our work and presents an auto-regressive denoising strategy for pretraining. While DPOT has shown better transferability to unseen PDE tasks than MPP, it shares the same limitations of focusing on 2D problems for pretraining and requiring large amount of data and compute (8 A800 GPUs for 500,000 steps).
@@
A final work that is closely related to ours is ORCA [16], which proposes a general workflow for adapting pretrained language/vision transformers to non-text, non-vision inputs. While ORCA uses PDEBench in its evaluation, it is not tailored to PDE solving and requires adapting a separate model for every dataset. The resulting models are not grid- or resolution-invariant, which are key properties of neural operators and achieved by UPS. Moreover, by learning from multiple PDEs and sharing knowledge across families, our method obtains significantly better empirical results than ORCA.
@@ 3 Methodology
Our goal is to train unified neural operators for spatiotemporal PDEs with varying domain dimensions, coefficients, initial and boundary conditions. These PDEs could model a range quantities that evolve over time, from scalars (e.g., pressure, density) to vectors (e.g., velocity). To achieve this, we propose UPS, which consists of a unified way to represent the PDE and a LLM-based network to model them.
@@ 3.1 Unified Data Representation
$\small\begin{split}\frac{d\bm{u}(t,\bm{x})}{dt}&=L\left(\bm{u}(t,\bm{x}),\frac{\partial\bm{u}(t,\bm{x})}{\partial\bm{x}},\frac{\partial\bm{u}(t,\bm{x})}{\partial\bm{x}^{2}},\cdots\right)\\ u(0,\bm{x})&=u_{0}(\bm{x})\;\;\;\;B\left(\bm{u}(t,\bm{y})\right)=0\end{split}$ (1)
@@
where $\bm{x}\in\Omega\subset\mathbb{R}^{d}$ is the spatial variable, $\bm{u}:[0,T]\times\Omega\to\mathbb{R}^{d_{u}}$ is a time-varying function defined over the domain $\Omega$ for finite time $T$. Here, $L$ is a (possibly non-linear) operator which acts on $\bm{u}$ and multiple partial derivatives of $\bm{u}$ w.r.t the spatial variable $\bm{x}$. $\bm{u}_{0}(\bm{x}):\Omega\to\mathbb{R}^{d_{u}}$ denotes PDE’s initial condition, and the operator $B$ defines the boundary condition where $\bm{y}\in\partial\Omega$ is a point on domain’s boundary. PDE families in this form include Navier-Stokes equations, Reaction-Diffusion equations, Burgers equations, and many others that describe phenomena like fluid dynamics and heat flow over time. They also constitute most PDE benchmarks in the field of machine learning [10, 37, 38].
@@
Consider a set of $S$ spatiotemporal PDEs $\{\bm{u}^{s}\}_{s=1}^{S}$. Here, each $\bm{u}^{s}=\{\bm{u}_{t}^{s}(\bm{x})\}_{t=1}^{T_{s}}$ is solution to a PDE of the form defined in Equation 1 such that for all $t\in[T_{s}]$, we have $\bm{u}^{s}_{t}(\bm{x})\in\mathbb{R}^{d^{s}_{u}}$ and $\bm{x}\in\Omega^{s}\subset\mathbb{R}^{d^{s}}$. For each $\bm{u}_{s}^{t}$, we assume that we have an $n$-point discretization of the functions $\{\bm{u}^{s}_{t}\}_{t=1}^{T_{s}}$ at points $W_{n}^{s}=\{x^{s}_{1},x^{s}_{2},\cdots,x^{s}_{n}\}$, where each $x^{s}_{i}\in\mathbb{R}^{d^{s}}$. That is, for each PDE $s\in S$ and $t\in T_{s}$, we have the realization of the function $\bm{u}^{s}_{t}$ on a grid with each dimension divided into $n$ parts. We assume that $n$ is constant across PDE families. Denote the set of $N$ physical quantities considered for each PDE as $V=\{v_{1},v_{2},\cdots v_{N}\}$. Our goal is to learn an operator ${\mathcal{G}}_{\theta}$ which, for a given PDE $s$, predicts the state of the PDE at time $t+1$ based on its state at time $t\in[T_{s}]$. That is:
$\bm{u}^{s}_{t+1}(\bm{x})={\mathcal{G}}_{\theta}(\bm{u}^{s}_{t}(\bm{x}))$
We thus need a unified representation for the inputs so a model can handle different quantities at once.
@@
Unifying Dimension Let $d=\max_{s\in S}d^{s}$. We want to represent all datasets in $\mathbb{R}^{d}$. Thus, for PDEs with $d^{s}<d$, the final $d-d^{s}$ coordinates of $x_{i}^{s}\in W_{n}^{s}$ are set to zero. In this work, we mainly consider PDEs defined over one- and two-dimensional domains, i.e., $d^{s}\in\{1,2\}\ \forall s\in S$. Hence, for PDEs with $d^{s}=1$, the point $x\in\Omega^{s}$ is represented as $(x,0)$. Note that our methodology to unify the total number of dimensions in the PDE is general and can be adapted to PDEs defined in higher-dimensional domains as well. In the following, we will denote $\bm{u}_{t}^{s}(\bm{x})$ as the value of the the function $\bm{u}_{t}^{s}$ on all the points in $W_{n}^{s}$, unless stated otherwise.
@@
Following the above procedure, we lift every PDE to a unified space such that $\bm{u}^{s}\in\mathbb{R}^{T_{s}\times N\times n^{d}}\ \forall s\in S$. To obtain the datasets for forward prediction, we generate input-output pairs via autoregressive teacher-forcing: for each time step $t\in[T_{s}]$, we use $\bm{u}^{s}_{t}$ to predict $\bm{u}^{s}_{t+1}$, yielding $T_{s}-1$ pairs of data from a single trajectory. We also append the coordinates of each $x_{i}^{s}\in W^{s}_{n}$ to the input and maintain an output mask for each sample, masking out the zero-padded dimensions when computing the loss.

### v4

@@ 1 Introduction
Partial Differential Equations (PDEs) play a pivotal role in modeling and understanding real-world phenomena, such as fluid dynamics and heat transfer. Although there exists a rich body of classical PDE solvers (Boyd, 2001; LeVeque, 2007; Moukalled et al., 2016) that are effective and mathematically proven, these solvers often incur substantial computational costs when used in practice, as they need to be re-run every time a coefficient or boundary condition changes. This motivates the development of neural operators (Li et al., 2020a; Chen & Chen, 1995; Lu et al., 2019), which use neural networks to approximate a solution map for a PDE family and can generalize to different initial/boundary conditions or coefficients. While existing neural operators (Lippe et al., 2023; Hao et al., 2023a; Marwah et al., 2023) have demonstrated strong performance on various practical benchmarks (Takamoto et al., 2022; Gupta & Brandstetter, 2022), most of them are designed to work with a single PDE family. Training a separate model for each PDE family remains costly.
@@
Several recent works, such as Subramanian et al. (2023), MPP (McCabe et al., 2023), and DPOT [^1] This work was done at the same time as ours. (Hao et al., 2024), have taken initial steps towards developing foundation models for PDE solving, learning unified operators that transfer across PDE families. These models are pretrained from scratch using extensive amounts of data and compute. For example, MPP is trained with over 80,000 PDE trajectories on 8 NVIDIA H100 GPUs for 200,000 steps. Despite the development costs, the resulting models are limited in generalization ability—all existing unified models focus on pretraining with 2D PDEs. Finally, as these models are developed using only PDE trajectories, they do not leverage meta information that could help distinguish between various PDE families, such as the name of the family and the coefficients.
@@
Figure 1: To adapt pretrained LLMs for PDE solving, UPS first transforms PDE of different dimensions, channels, and resolutions into a unified representation (left panel). Then, the data is processed with a unified architecture that integrates FNO layers, PDE metadata, and LLMs (right panel). The architecture is trained in two stages. In stage 1, we pretrain the embedding network using a joint loss that simultaneously optimizes (i) the distribution similarity between PDE features and text embeddings to align the modalities, and (ii) the prediction performance of extracted PDE features. In stage 2, we fine-tune the entire model on a dataset that combines multiple families of spatiotemporal PDEs with varying domain dimensions, initial/boundary conditions, and coefficients. UPS achieves competitive results with significantly better sample-efficiency than existing methods.
@@
We present Unified PDE Solvers (UPS), which learns unified neural operators for complex time-dependent PDEs with improved efficiency and generalization ability. Unlike existing efforts that train models from scratch, we propose a novel method to adapt pretrained Large Language Models (LLMs) to PDE solving. This is inspired by the line of work that repurposes LLMs for scientific domains like mathematics (Lewkowycz et al., 2022), computational biology (Shen et al., 2023; Vinod et al., 2023; Joachimiak et al., 2023), and chemistry (Bran et al., 2023; Shen et al., 2024). These works not only show how LLMs utilize both text and non-text information to solve scientific problems and transfer effectively to unseen tasks, but also provide strong evidence that the general-purpose pretraining and inductive biases of LLMs could substantially reduce the sample complexity needed for adaptation.
@@
    We propose a unified data representation scheme to align PDEs with varying dimensions and physical quantities into the same feature space. Given the space and time discretization $\bm{u}=\{\bm{u}_{t}(\bm{x})\}_{t=0}^{T}$, where $\bm{x}\in\mathbb{R}^{d}$ is the spatial variable and $\bm{u}_{t}(\bm{x})$ is the state variable, UPS homogenizes $\bm{u}_{t}(\bm{x})$ from diverse PDEs into a shared “superspace” $\mathbb{R}^{N\times n^{d_{\max}}}$, where $d_{\max}$ is the maximum dimension of $\bm{x}$ among all PDEs considered, $N$ is the superset of the physical quantities, and $n$ is the resolution. This framework of embedding lower-dimensional PDEs in a higher dimension enables UPS to model cross-dimensional PDEs simultaneously and distinguishes us from all existing unified operators, which do not consider low dimensional PDEs in 1D.
@@
    We employ a unified network architecture to predict $\bm{u}_{t+1}(\bm{x})$ based on $\bm{u}_{t}(\bm{x})$. To leverage pretrained LLMs, we design a three-way architecture that consists of (i) a FNO (Li et al., 2020a) based embedding network to convert PDE data into resolution-invariant, sequential features; (ii) an LLM body to process the PDE features and the text embeddings of the PDE metadata; and (iii) a prediction network to generate the final output. Inspired by previous cross-modality adaptation work (Shen et al., 2023), we employ a two-stage align-then-refine process for model training. However, we improve the align stage by using a joint loss that adds feature extraction on top of alignment to pretrain the embedding network. We improve the refine stage by fine-tuning on a collection of PDE tasks rather than a single task. Our enhanced workflow outperforms naive transfer and previous cross-modal approaches, reducing both the data and compute needed for training.
By design, UPS can handle diverse PDE families, data dimensions, channels, and resolutions. More crucially, by warm-starting with pretrained LLM weights and applying explicit alignment, UPS strikes a balance between effectiveness and efficiency—it achieves state-of-the-art performance across 9 datasets from PDEBench (Takamoto et al., 2022) (7 in-distribution, 2 out-of-distribution), using about 20,000 training trajectories, a single A6000, 60,000 train steps, and under 100 GPU hours. This means that we achieve better results than existing unified models using 4 times less data and 26 times less compute.
Beyond prediction accuracy, we confirm that UPS preserves key properties of neural operators, such as grid- and resolution-invariance. We also show that UPS is compatible with a variety of LLM backbones, including RoBERTa (Liu et al., 2019), T5 (Raffel et al., 2020), and CLIP (Radford et al., 2021), and demonstrates better performance when scaled to larger backbones. We believe that the model-agnostic design of UPS offers a systematic approach to harnessing the advancements in LLMs for PDE solving, and it takes a further step towards building generalized foundation models for more complex physical systems efficiently. Code is available at https://github.com/sjunhongshen/UnifiedPDESolvers.
@@ 2 Related Work
Recent years has seen a variety of neural-network-based methods for approximating PDE solutions. Hybrid solvers (Hsieh et al., 2019; Bar-Sinai et al., 2019; Kochkov et al., 2021) apply classical solvers like finite element/volumn methods (LeVeque, 2007; Moukalled et al., 2016) to a low-resolution grid and use neural networks to predict the correction terms. Others directly approximate the PDE solutions with neural networks (Sirignano, 2017; Raissi et al., 2019; Khoo et al., 2021; Shen et al., 2022), using variational losses (Yu et al., 2018) or physical constraints defined by the PDE (Raissi et al., 2019; Bruna et al., 2024). Being mostly equation-specific, these methods can solve one PDE at a time. The learned models do not apply to other PDEs in the same family, let alone other families.
@@
A more general approach involves learning neural operators (Lu et al., 2019; Li et al., 2020a; b) which approximate an infinite-dimensional operator between two functional spaces. For time-dependent PDEs, a neural operator maps the current state of a PDE to the next state, with quantities like initial conditions provided as input. Neural operators can be implemented using any architecture. For example, Fourier neural operator (FNO) (Li et al., 2020a) uses convolution-based integral kernels evaluated in the Fourier space. Other works also use transformer models (Cao, 2021; Li et al., 2022; Hao et al., 2023a) or U-Net (Lippe et al., 2023). Learning neural operators enables solving an entire family of PDE and they can easily adapt to new parameterizations of a PDE without fine-tuning. However, the learned operators cannot extend to different PDE families.
@@
To facilitate operator transfer across PDE families, two recent works develop large pretrained models for multiple physical systems: Subramanian et al. (2023) train FNOs on steady-state linear PDEs with periodic boundary conditions; McCabe et al. (2023) design a new transformer architecture based on the axial attention (Ho et al., 2020) and train it using various 2D non-linear, time-dependent PDEs. While these methods show that a unified operator can outperform single-family operators, they are limited in two aspects. First, existing unified methods consider mainly 2D PDEs for pretraining and evaluation. In contrast, UPS leverages a unified representation scheme to tackle both 1D and 2D PDEs. This method can be also extended to any $d$-dimensional systems in theory. Second, existing methods pretrain large models from scratch and necessitate extensive GPU resources and pretraining data, which can be prohibitive to collect for high-dimensional complex PDEs. However, by adapting from pretrained LLMs and closing the modality gap between text and PDE efficiently, UPS achieves competitive results using 4x less data and 26x less compute.
@@
Beyond the aforementioned works, DPOT (Hao et al., 2024) was developed concurrently with our work and presents an auto-regressive denoising strategy for pretraining. While DPOT has shown better transferability to unseen PDE tasks than MPP, it shares the same limitations of focusing on 2D problems for pretraining and requiring large amount of data and compute (8 A800 GPUs for 500,000 steps).
@@
A final work that is related to ours is ORCA (Shen et al., 2023), which proposes a general workflow for adapting pretrained language/vision transformers to non-text/vision inputs. While ORCA uses PDEBench in its evaluation, it is not tailored to PDE solving and requires adapting a separate model for every dataset. The resulting models are not grid- or resolution-invariant, which are key properties of neural operators and achieved by UPS. Moreover, by learning from multiple PDEs and sharing knowledge across families, UPS obtains significantly better empirical results than ORCA.
@@ 3 Methodology
Our goal is to train unified neural operators for spatiotemporal PDEs with varying domain dimensions, coefficients, initial and boundary conditions. These PDEs could model a range of quantities that evolve over time, from scalars (e.g., pressure, density) to vectors (e.g., velocity). To achieve this, we propose UPS, which consists of a unified way to represent the PDE and a LLM-based network to model them.
@@ 3.1 Unified Data Representation
$\begin{split}\frac{\partial\bm{u}(t,\bm{x})}{\partial t}&=L\left(\bm{u}(t,\bm{x}),\frac{\partial\bm{u}(t,\bm{x})}{\partial\bm{x}},\frac{\partial\bm{u}(t,\bm{x})}{\partial\bm{x}^{2}},\cdots\right)\\ u(0,\bm{x})&=u_{0}(\bm{x})\;\;\;\;B\left(\bm{u}(t,\bm{y})\right)=h(y)\end{split}$ (1)
@@
where $\bm{x}\in\Omega\subset\mathbb{R}^{d}$ is the spatial variable, $\bm{u}:[0,T]\times\Omega\to\mathbb{R}^{d_{u}}$ is a time-varying function defined over the domain $\Omega$ for finite time $T$. Here, $L$ is a (possibly non-linear) operator which acts on $\bm{u}$ and multiple partial derivatives of $\bm{u}$ w.r.t the spatial variable $\bm{x}$. $\bm{u}_{0}(\bm{x}):\Omega\to\mathbb{R}^{d_{u}}$ denotes PDE’s initial condition, and the operator $B$ defines the boundary condition where $\bm{y}\in\partial\Omega$ is a point on domain’s boundary, and $h:\partial\Omega\to\mathbb{R}^{d_{u}}$ defines the given function over the boundary [^2] Unless stated otherwise, we assume that the value of the function is 00 at the boundary, i.e, $h(y)=0$ for all $y\in\partial\Omega$.. PDE families in this form include Navier-Stokes equations, Reaction-Diffusion equations, Burgers equations, and many others that describe phenomena like fluid dynamics and heat flow over time. They also constitute most PDE benchmarks in the field of machine learning (Takamoto et al., 2022; Tu et al., 2022; Roberts et al., 2021).
@@
Consider a set of $S$ spatiotemporal PDEs $\{\bm{u}^{s}\}_{s=1}^{S}$. Here, each $\bm{u}^{s}=\{\bm{u}_{t}^{s}(\bm{x})\}_{t=1}^{T_{s}}$ is a solution to a PDE of the form defined in equation 1 such that for all $t\in[T_{s}]$, we have $\bm{u}^{s}_{t}(\bm{x})\in\mathbb{R}^{d^{s}_{u}}$ and $\bm{x}\in\Omega^{s}\subset\mathbb{R}^{d^{s}}$, where $d^{s}$ is the dimension of the PDE $s$. For each $\bm{u}_{t}^{s}$, we assume that we have an $n$-point discretization of the functions $\{\bm{u}^{s}_{t}\}_{t=1}^{T_{s}}$ at points $W_{n}^{s}=\{\bm{x}^{s}_{1},\bm{x}^{s}_{2},\cdots,\bm{x}^{s}_{n^{d^{s}}}\}$, where each $\bm{x}^{s}_{i}\in\mathbb{R}^{d^{s}}$. That is, for each PDE $s\in S$ and $t\in T_{s}$, we have the realization of the function $\bm{u}^{s}_{t}$ on a grid with each dimension divided into $n$ parts, thus giving rise to $n^{d^{s}}$ points in the set. We assume that $n$ is constant across PDE families. We note that this value $n$ is a hyperparameter, and ideally should be the minimum number of points that work well for all the PDEs considered, for example, low-viscosity Navier-Stokes may require more discretization points compared to the high viscosity counterparts. Denote the set of $N$ physical quantities considered for each PDE as $V=\{v_{1},v_{2},\cdots v_{N}\}$. Our goal is to learn an operator ${\mathcal{G}}_{\theta}$ which, for a given PDE $s$, predicts the state of the PDE at time $t+1$ based on its state at time $t\in[T_{s}]$, i.e., $\hat{\bm{u}}^{s}_{t+1}(\bm{x})={\mathcal{G}}_{\theta}(\bm{u}^{s}_{t}(\bm{x}))$. We thus need a unified representation for the inputs so a model can handle different quantities at once.
@@
Unifying Dimension Let $d^{s}$ denote the dimension of the PDE $s$ and $d=\max_{s\in S}d^{s}$. We want to represent all datasets in $\mathbb{R}^{d}$. Thus, for PDEs with $d^{s}<d$, the final $d-d^{s}$ coordinates of $\bm{x}_{i}^{s}\in W_{n}^{s}$ are set to zero. In this work, we mainly consider PDEs defined over one- and two-dimensional domains, i.e., $d^{s}\in\{1,2\}\ \forall s\in S$. Hence, for PDEs with $d^{s}=1$, the point $\bm{x}\in\Omega^{s}$ is represented with the 2D-coordinate $(x,0)$. Note that our methodology to unify the total number of dimensions in the PDE is general and can be adapted to PDEs defined in higher-dimensional domains as well. In the following, we will denote $\bm{u}_{t}^{s}(\bm{x})$ as the value of the function $\bm{u}_{t}^{s}$ on all the points in $W_{n}^{s}$, unless stated otherwise.
@@
With the above procedure, we lift every PDE to a unified space so $\bm{u}^{s}\in\mathbb{R}^{T_{s}\times N\times n^{d}}\ \forall s\in S$. To obtain the datasets for forward prediction, we generate input-output pairs via autoregressive teacher-forcing: for each time step $t\in[T_{s}]$, we use $\bm{u}^{s}_{t}$ to predict $\hat{\bm{u}}^{s}_{t+1}$, yielding $T_{s}-1$ pairs of data from a single trajectory. We append the coordinates of each $\bm{x}_{i}^{s}\in W^{s}_{n}$ to the input and maintain an output mask to mask out the zero-padded dimensions when computing the loss.